I have to find the exact value for this expression: sin 260 degrees cos 20 degrees - cos 260 degrees sin 20 degrees
help?

Answer by Boreal(15235)

This is of the form sin a*cos b-cos a*sin b
That is sin (a-b)=sin (260deg-20deg)=sin (240 deg)
That is the same as -sin (60 deg)=- sqrt(3)/2